OpenStudy (anonymous):

Factorize (2g)^2 - 18

OpenStudy (anonymous):

2(2g^2-9)

OpenStudy (anonymous):

using this identity a^2-b^2=(a+b)(a-b) is it possible to solve it

OpenStudy (anonymous):

Yes, you can \[18=(3\sqrt{2})^2\] So we have \[(2g)^2-(3\sqrt{2})^2\] apply the identity as required!

OpenStudy (anonymous):

or you can apply it after factoring out a 2 \[(2g)^2-18=4g^2-18=2(2g^2-9)=2((\sqrt{2}g)^2-3^2)\]